Find the derivative of the following:
y=(8-2x^2)(4-x)
1
Expert's answer
2012-11-19T08:02:58-0500
y'=(8-2x^2)'*(4-x)+(8-2x^2)*(4-x)'=-4x(4-x)-(8-2x^2)=-16x+4x^2-8+2x^2=6x^2-16x-8